# On the Vulnerability of Concept Erasure in Diffusion Models

### v2

@@ Abstract
The proliferation of text-to-image diffusion models has raised significant privacy and security concerns, particularly regarding the generation of copyrighted or harmful images. In response, several concept erasure (defense) methods have been developed to prevent the generation of unwanted content through post-hoc finetuning. On the other hand, concept restoration (attack) methods seek to recover supposedly erased concepts via adversarially crafted prompts. However, all existing restoration methods only succeed in the highly restrictive scenario of finding adversarial prompts tailed to some fixed seed. To address this, we introduce RECORD, a novel tangential- coordinate-descent-based restoration algorithm that finds adversarial prompts to recover erased concepts independently of the pseudo-random number seed. Our extensive experiments demonstrate RECORD consistently outperforms the current restoration methods by up to 17.8 times in this setting. Our findings further reveal the susceptibility of unlearned models to restoration attacks, providing crucial insights into the behavior of unlearned models under the influence of adversarial prompts. The code for RECORD is available at [available upon publication].
@@
[…]
figurea) Examples of restored concepts from models unlearned on van Gogh painting style. b) The update schematic of RECORD, which uses a linear gradient approximation to obtain a small set of candidate tokens, and then updates the prompt with respect to their exact gradients.
@@ 1 Introduction
Text-to-Image Diffusion models have recently garnered significant attention for their ability to generate high-quality images from natural language inputs [36, 30]. However, because these models are trained on vast and diverse datasets that may contain harmful or undesirable content, their proliferation raises substantial ethical and safety concerns, particularly over the generation of private, copyrighted and harmful content [3, 35]. Pre-filtering undesired images from the training dataset is often considered impractical due to the sheer size of these datasets, as well as the necessity to re-train the model from scratch. Consequently, much research has pursued post-hoc approaches aiming to remove the influence of undesired content from trained models via low-cost finetuning, while preserving the generation quality of other non-erased concepts [10, 43, 42, 45, 20, 25, 8, 11, 46, 12, 18]. This is commonly referred to as concept erasure, a subfield of machine unlearning [19].
@@
However, it is well-known that neural networks are susceptible to adversarial attacks: small perturbations to an input can induce a well-trained model to produce any pre-determined outputs without altering the model itself [21, 5, 1]. This vulnerability raises similar concerns in the context of concept erasure. Indeed, recent studies have largely demonstrated the feasibility of eliciting unlearned models to re-generated the erased concepts via white-box optimization-based attacks [3, 47]. Although the reliance on a white-box scenario limits the practicality of these attacks by external adversaries on closed-sourced models, these methods are instrumental in auditing the robustness in concept erasure research. We refer to this class of attack methods as concept restoration.
The existing concept restoration methods rely on projecting the discrete text prompts into a continuous and differentiable space to enable gradient-based optimization [3, 47]. However, projection-based adversarial attacks have been shown to perform poorly on large language models (LLMs) [2]. Consequently, all existing restoration methods can only identify adversarial prompts tailored to a specific seed, which an outside adversary likely has no control over. In particular, our experiments demonstrate that these methods almost completely fail when evaluating their optimized prompts with different random seeds. In this paper, we propose RECORD (Restoring Erased Concepts via Coordinate Descent), a white-box tangential- coordinate-descent algorithm that employs a two-stage optimization scheme that negates the need for projection. To the best of our knowledge, RECORD is the first concept restoration method to consistently find non-seed-specific adversarial prompts that can restore a whole distribution of images of an erased concept, achieving a 17.8-fold improvement over the existing state-of-the-art restoration methods. Examples of restored images are represented in Table 1
@@
Our investigations also reveal crucial insights into the fundamental vulnerability of the unlearned models. Notably, for any arbitrary initialization in the prompt embedding space, our findings suggest there always exist nearby embeddings that can trigger the generation of erased concepts, suggesting possible adversarial prompts exist almost everywhere in the prompt embedding space. Furthermore, embeddings initialized close to the embeddings of exact descriptions of the erased concept tend to move away from those embeddings during optimization. This suggests most existing unlearning methods only suppress the generation of the erased concept when using its exact description or its semantically similar versions.
@@ 2.1 Text-to-Image Diffusion Models
Diffusion Models are a class of generative model that generate images from text by learning to reverse the forward diffusion process. Starting with Gaussian noise $x_{T}\sim\mathcal{N}(0,\mathbf{1})$, a trained denoiser, commonly a U-Net [31] or Vision Transformer [6], iteratively denoises $x_{T}$ over the interval $t\in[0,T]$ until a clear image $x_{0}$ is reached. By conditioning on prompt embeddings $c=\mathcal{T}(y)$, where $y$ is some natural language prompt, text-to-image generation is achieved. $\mathcal{T}$ is a pre-trained text encoder, commonly CLIP [28] or BLIP [24]. In practice, many popular diffusion models, such as Stable Diffusion [30], perform the denoising in a latent space $z_{t}=\mathcal{E}(x_{t})$. The denoiser $\epsilon_{\theta}$ parameterized by $\theta$ can then be trained by optimizing
@@
where $z_{t}$ is obtained by applying the forward diffusion process to the clean latent $z_{0}$ using the Gaussian noise $\epsilon$.
@@
[10] [45] [20] [25] [11] [46]
P4D [3] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
UD [47] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 1: Generated images of erased concepts (van Gogh paintings) using prompt attacks. Each image column of the same concept is generated using the same latent initialization.
@@ 2.2 Prompt Tuning
Manipulating prompts to elicit specific behaviors from language models, also known as prompt tuning, is an important topic in Natural Language Processing research. [7] introduced HotFlip, generating adversarial examples through minimal character-level flips guided by gradients. Extending this, [39] developed Universal Adversarial Triggers—input-agnostic token sequences optimized by using first order Taylor-expansion around the current token to select candidate tokens for exact evaluation. [34] presented AutoPrompt, designed to automatically generate prompts for various use cases. Addressing the lack of fluency in these prompts, [33] introduced FluentPrompt, incorporating fluency constraints and using Langevin Dynamics combined with Projected Stochastic Gradient Descent, where projection is done onto the set of token embeddings. [41] developed PEZ, an algorithm inspired by FluentPrompt, allowing for prompt creation in both text-to-text and text-to-image applications. In text-to-image models, [9] applied Textual Inversion, learning “pseudo-words” in the embedding space to capture specific visual concepts. Further advancements include GBDA [13], enabling gradient-based optimization over token distributions using the Gumbel-Softmax reparametrizaton [16] to stay on the probability simplex, and ARCA [17], optimizing discrete prompts via an improvement to AutoPrompt. ARCA will inspire our method.
@@ 2.3 Concept Restoration
Recent methods designed to restore erased concepts from unlearned models often leverage advanced optimization techniques similar to prompt tuning. Concept Inversion (CI) [27] introduces a new token with a learnable embedding to represent the erased concept, which is learned to minimizes the reconstruction loss during denoising. This is a direct application of Textual Inversion [9] from prompt tuning to the concept restoration paradigm. Prompting Unlearned Diffusion Models (PUND) [14] enhances this approach by iteratively erasing and searching for the concept by also updating model parameters, improving transferability across models.
Other methods focus on discrete token optimization. UnlearnDiffAtk (UD) [47] performs optimization over token distributions, similar to GBDA [13], but utilizes projection onto the probability simplex instead of the Gumbel-Softmax reparameterization. Prompting4Debugging (P4D) [3] optimizes prompts in the embedding space and projects onto discrete embeddings, akin to the approach used in PEZ [41]. Additionally, Ring-A-Bell [38] constructs an empirical representation of the erased concept by averaging embedding differences from prompts with and without the concept, then employs a genetic algorithm to optimize the prompt.
3 Method
3.1 Motivation: Vulnerability of Machine Unlearning
Vulnerability Verifying whether a model has truly unlearned a concept is challenging. To assess the effectiveness of the unlearning process, we consider an unlearned denoiser $\epsilon_{\theta^{\prime}}$ to be robust if it consistently fails to generate the erased content and produce images significantly different from those generated by the original model $\epsilon_{\theta}$ when subjected to any adversarial prompt and any latent initialization. Therefore, this work focuses on measuring the degree to which the unlearned model has diverged from the original model concerning the erased content. To achieve this, we propose a loss function similar to [3]
@@ 3.1 Motivation: Vulnerability of Machine Unlearning
This formulation is similarly applicable to erasure methods which unlearns the text encoder $\mathcal{T}$.
Some literature recommends the use of a Gaussian noise $\epsilon$ in place of the original model $\epsilon_{\theta^{\prime}}$, citing the reduced computational costs [47]. Empirically we observe the loss function from equation (1) improves the effectiveness of concept restoration attacks, as the predictions of the original model $\epsilon_{\theta}$ act as a more informative surrogate than the Gaussian noise. We refer the readers to appendix A for more detailed discussions and experimental results.
This paper considers two types of restoration attacks to assess the vulnerability of unlearned models:
@@
  • Prompt embedding attacks: In this setting, concept restoration is achieved by directly perturbing the prompt embedding $c$ to minimize the loss function defined in equation (1). With the prompt embedding space being continuous and differentiable, finding non-seed-specific adversarial prompts poses an easier task. However, precise inversions from embeddings back to prompts are not guaranteed to exist, and embedding attacks are less practicable in most scenarios. Here we use these attacks as ablation studies in section 4.3 for analyzing concept erasure vulnerability under small continuous perturbations, and in appendix A for examining different loss functions.
@@
  • Prompt attacks: Directly perturbing prompt tokens to restore concepts is significantly more challenging due to their discrete and non-differentiable nature. Consequently, existing restoration methods primarily focus on pseudo-random number seed-specific adversarial prompts and struggle to find general adversarial prompts effective across all seeds. To overcome this limitation, we introduce RECORD (Restoring Erased Concepts via Coordinate Descent) for carrying out robust non-seed-specific concept restoration attacks.
3.2 RECORD
The tangential- coordinate-descent-based two-stage approach employed by RECORD iteratively optimizes the prompt one token at a time while fixing all other tokens in the prompt sequence. The algorithm first initializes a random token sequence $y=[y_{1},\ldots,y_{S}]$ of length $S$, and iteratively perform $N$ passes over $y$. In each pass, a random permutation of the token positions is generated to mitigate positional bias during updates. For each position $s$ in the permuted sequence, the algorithm samples a batch of $K$ images $z_{0}^{[n]}$ and corresponding timesteps $t^{[n,s]}$. Candidate tokens $v$ for position $s$ are then selected by sampling $J$ random tokens. Since computing the exact expectation over all latents and timesteps is intractable, we approximate the $L(y)$ from equation (1) as:
@@ 3.2 RECORD
$\hat{L}(y,\mathcal{Z})={}\sum_{(z_{t},t)\in\mathcal{Z}}\Bigl{\|}\epsilon_{\theta^{\prime}}\left(z_{t},t,\mathcal{T}(y)\right)-\epsilon_{\theta}\left(z_{t},t,\mathcal{T}\left(y_{\text{target}}\right)\right)\Bigr{\|}_{2}^{2},$ (2)
where $\mathcal{Z}$ is a sampled batch of noised images and their corresponding timesteps. The gradient $g$ of $\hat{L}$ with respect to the candidate token embeddings can thus be computed as:
$g_{j}=\nabla{c_{j}}\hat{L}(y({c_{j}},s),\mathcal{Z}(n,s)).$

### v3

# Rethinking the Vulnerability of Concept Erasure and a New Method

@@ Abstract
The proliferation of text-to-image diffusion models has raised significant privacy and security concerns, particularly regarding the generation of copyrighted or harmful images. In response, concept erasure (defense) methods have been developed to "unlearn" specific concepts through post-hoc finetuning. However, recent concept restoration (attack) methods have demonstrated that these supposedly erased concepts can be recovered using adversarially crafted prompts, revealing a critical vulnerability in current defense mechanisms. In this work, we first investigate the fundamental sources of adversarial vulnerability and reveal that vulnerabilities are pervasive in the prompt embedding space of concept-erased models, a characteristic inherited from the original pre-unlearned model. Furthermore, we introduce RECORD, a novel tangential- coordinate-descent-based restoration algorithm that consistently outperforms existing restoration methods by up to 17.8 times. We conduct extensive experiments to assess its compute-performance tradeoff and propose acceleration strategies. The code for RECORD is available at [available upon publication].
@@
Figure 1: a) Examples images from models unlearned on van Gogh painting style. b) The update schematic of RECORD, which uses a linear gradient approximation to obtain a small set of candidate tokens, and then updates the prompt with respect to the exact evaluation of the loss function.
@@ 1 Introduction
Text-to-Image Diffusion models have recently garnered significant attention for their ability to generate high-quality images from natural language inputs Song et al. (2020); Rombach et al. (2022). However, because these models are trained on vast and diverse datasets that may contain harmful or undesirable content, their proliferation raises substantial ethical and safety concerns, particularly over the generation of copyrighted and harmful content Chin et al. (2023); Somepalli et al. (2022). Pre-filtering undesired images from the training dataset is often considered impractical due to the sheer size of these datasets, as well as the cost of re-training models from scratch. Consequently, much research has pursued post-hoc approaches aiming to remove the undesired content from trained models via low-cost finetuning, while preserving the generation quality of other non-erased concepts Gandikota et al. (2023a); Wu et al. (2024); Wu & Harandi (2024); Zhang et al. (2024a); Kumari et al. (2023); Lyu et al. (2024); Fan et al. (2023); Gandikota et al. (2023b); Zhang et al. (2024b); Gong et al. (2024); Kim et al. (2024); Zhang et al. (2025); Gao et al. (2024); Srivatsan et al. (2025). This is commonly referred to as concept erasure, a subfield of machine unlearning Kim & Qi (2025).
@@
However, it is well-known that neural networks are susceptible to adversarial attacks: small perturbations to an input can induce a well-trained model to produce any pre-determined outputs without altering the model itself Kurakin et al. (2016); Dong et al. (2018); Yang et al. (2024); Beerens & Higham (2024). This vulnerability raises similar concerns in the context of concept erasure. Indeed, recent studies have largely demonstrated the feasibility of eliciting unlearned models to re-generate the erased concepts via white-box optimization-based attacks Chin et al. (2023); Zhang et al. (2023). We refer to this class of attack methods as concept restoration.
@@
While the success of concept restoration methods demonstrates the vulnerability of concept-erased models, the underlying cause of this persistent susceptibility to adversarial attacks is not yet fully understood Lu et al. (2025). To this end, our investigations reveal crucial insights into the fundamental vulnerability of unlearned models. Notably, prompt embeddings initialized in different regions of the embedding space, our findings suggest there often exist nearby embeddings that can restore the erased concepts. This suggests adversarial embeddings are pervasive in the prompt embedding space and can be exploited by the existing restoration algorithms. Furthermore, for certain types of erasure methods, embeddings initialized near the exact descriptions of the erased concept tend to diverge from those embeddings during optimization. This suggests that most existing unlearning methods only suppress the generation of the erased concept under prompts embedded near the specific prompt embeddings corresponding to the erased concept.
Stemming from these findings, we notice existing concept restoration methods rely on projecting the discrete text prompts into a continuous and differentiable space to enable gradient-based optimization Chin et al. (2023); Zhang et al. (2023). However, recent studies have demonstrated that projection-based adversarial attacks generally underperform in comparison to coordinate-descend-based approaches Carlini et al. (2023); Zou et al. (2023); Jones et al. (2023) in language model adversarial attacks. This motivates our investigation of similar approaches in the field of concept restoration. Therefore, we further propose RECORD (Restoring Erased Concepts via Coordinate Descent), a white-box tangential- coordinate-descent algorithm employing a two-stage optimization scheme to negate the need for projection (Figure 1). Our extensive experiments demonstrated that RECORD consistently achieve superior performance by up to 17.8-fold over the existing state-of-the-art restoration methods. Examples of the restored images are presented in Table 1.
The contributions of this paper are as follows:
  • We explore why the majority of the current concept erasure methods are largely susceptible to concept restoration attacks.
  • We extend the existing concept restoration attack methods by introducing RECORD, a coordinate descent approach motivated by similar successes on language model adversarial attacks.
  • We conduct extensive ablation studies on RECORD, carefully assessing the effect of each hyperparameter and revealing its highly flexible compute-performance tradeoffs.
@@ 2.1 Text-to-Image Diffusion Models
Diffusion Models are a class of generative model that generate images from text by learning to reverse the forward diffusion process. Starting with Gaussian noise $x_{T}\sim\mathcal{N}(0,\mathbf{1})$, a trained denoiser, commonly a U-Net Ronneberger et al. (2015) or Vision Transformer Dosovitskiy et al. (2020), iteratively denoises $x_{T}$ over the interval $t\in[0,T]$ until a clear image $x_{0}$ is reached. By conditioning on prompt embeddings $c=\mathcal{T}(y)$, where $y$ is some natural language prompt, text-to-image generation is achieved. $\mathcal{T}$ is a pre-trained text encoder, commonly CLIP Radford et al. (2021) or BLIP Li et al. (2022). Latent diffusion models, such as Stable Diffusion Rombach et al. (2022), perform the denoising in latent space $z_{t}=\mathcal{E}(x_{t})$ and the denoiser $\epsilon_{\theta}$ is trained with the following objective
@@
where $z_{t}$ is obtained from the forward diffusion process to the clean latent $z_{0}$ with Gaussian noise $\epsilon$.
@@
2023a 2024a 2023 2024 2023b 2024b
P4D 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
UD 2023 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 1: Example images of erased concepts (van Gogh painting style) using token-level attacks. Each image column of the same concept is generated using the same latent initialization.
@@ 2.2 Prompt Tuning
Manipulating prompts to elicit specific behaviors from language models, also known as prompt tuning, is an important topic in Natural Language Processing research. Ebrahimi et al. (2018) introduced HotFlip, generating adversarial examples through minimal character-level flips guided by gradients. Extending this, Wallace et al. (2021) developed Universal Adversarial Triggers—input-agnostic token sequences optimized by using first order Taylor-expansion around the current token to select candidate tokens for exact evaluation. Shin et al. (2020) presented AutoPrompt, designed to automatically generate prompts for various use cases. Addressing the lack of fluency in these prompts, Shi et al. (2022) introduced FluentPrompt, incorporating fluency constraints and using Langevin Dynamics combined with Projected Stochastic Gradient Descent, where projection is done onto the set of token embeddings. Wen et al. (2023) developed PEZ, an algorithm inspired by FluentPrompt, allowing for prompt creation in both text-to-text and text-to-image applications. In text-to-image models, Gal et al. (2022) applied Textual Inversion, learning "pseudo-words" in the embedding space to capture specific visual concepts. Further advancements include GBDA Guo et al. (2021), enabling gradient-based optimization over token distributions using the Gumbel-Softmax reparametrizaton Jang et al. (2017) to stay on the probability simplex, GCG Zou et al. (2023) and ARCA Jones et al. (2023), optimizing discrete prompts via an improvement to AutoPrompt. ARCA will inspire our method.
@@ 2.3 Concept Restoration
Recent methods for restoring erased concepts from unlearned models often leverage advanced optimization techniques similar to prompt tuning. Concept Inversion (CI) Pham et al. (2023) introduces a new token with learnable embedding to represent the erased concept, which is learned to minimizes the reconstruction loss during denoising. This is a direct application of Textual Inversion Gal et al. (2022) from prompt tuning to the concept restoration paradigm. Prompting Unlearned Diffusion Models (PUND) Han et al. (2024) enhances this approach by iteratively erasing and searching for the concept while also updating model parameters, improving transferability across models.
Other methods focus on discrete token optimization. UnlearnDiffAtk (UD) Zhang et al. (2023) performs optimization over token distributions, similar to GBDA Guo et al. (2021), but utilizes projection onto the probability simplex instead of the Gumbel-Softmax reparameterization. Prompting4Debugging (P4D) Chin et al. (2023) optimizes prompts in the embedding space and projects onto discrete embeddings, akin to the approach used in PEZ Wen et al. (2023). Additionally, Ring-A-Bell Tsai et al. (2023) constructs an empirical representation of the erased concept by averaging embedding differences from prompts with and without the concept, then employs a genetic algorithm to optimize the prompt.
3 Methods
3.1 Motivation
Verifying whether a model has truly unlearned a concept is challenging. To assess the effectiveness of the unlearning process, we consider an unlearned denoiser $\epsilon_{\theta^{\prime}}$ to be robust if it consistently fails to generate the erased content and produce images significantly different from those generated by the original model $\epsilon_{\theta}$ when subjected to any adversarial prompt and any latent initialization. Therefore, this work focuses on measuring the degree to which the unlearned model has diverged from the original model concerning the erased content. To achieve this, we propose a loss function similar to Chin et al. (2023)
@@ 3.1 Motivation
This formulation is similarly applicable to erasure methods which unlearns the text encoder $\mathcal{T}$. This paper considers two types of restoration attacks to assess the vulnerability of unlearned models:
@@
  • Embedding-level attacks: In this setting, concept restoration is achieved by directly perturbing the prompt embedding $c$ to minimize the loss function defined in Equation [1]. With the prompt embedding space being continuous and differentiable, finding adversarial prompts poses an easier task. However, precise inversions from embeddings back to prompts are not guaranteed to exist, making embedding-level attacks less practicable and realistic in most circumstances.
@@
  • Token-level attacks: Directly perturbing prompt tokens to restore concepts is significantly more challenging due to their discrete and non-differentiable nature. To overcome this limitation, we introduce RECORD for carrying out robust concept restoration attacks.
Figure 2: Behavior of the text embeddings during embedding-level attacks on models unlearned with ESD and AdvUnlearn. a) Isomap projection of the optimization trajectories in the prompt embedding space $\mathbb{R}^{T\times 77\times 768}$ down to $\mathbb{R}^{T\times 2}$. 2000 trajectories shown, each $T=10$ steps long. Dots / crosses denote the starting point. The erased concept can be generated at the end of each trajectory. b), c), d) present the cosine similarity histogram, computed in $\mathbb{R}^{77\times 768}$, between the initial, optimized, and reference target embeddings.
3.2 Embedding-Level Attacks: A Peek into the Origin of Model Vulnerability
Embedding-level attacks for restoring the erased concept can be easily carried out in a naive approach of directly optimizing on the prompt embedding $c$ with a fixed learning rate. Since computing the exact expectation over all latents and timesteps is intractable, we approximate the $L(y)$ from Equation [1] as:
@@ 3.2 Embedding-Level Attacks: A Peek into the Origin of Model Vulnerability
$\hat{L}(c,\mathcal{Z})={}\sum_{(z_{t},t)\in\mathcal{Z}}\Bigl\|\epsilon_{\theta^{\prime}}\left(z_{t},t,c\right)-\epsilon_{\theta}\left(z_{t},t,\mathcal{T}\left(y_{\text{target}}\right)\right)\Bigr\|_{2}^{2},$ (2)
where $\mathcal{Z}$ is a sampled batch of noised images and their corresponding timesteps. For embedding-level attacks, we use a batch size of 16 images and NAdam for optimizing onprompt embedding $c$ with a fixed learning rate of 0.1. The full embedding-level attack results can be found in Appendix B.
To explore vulnerability, we visualize the 2D isomap Wang (2012) projections of prompt embedding optimization trajectories (Figure 2a), providing a visual intuition for their flow within the embedding space. We use an exact description of the erased concept as the reference target, e.g. "a painting in the style of van Gogh" for models unlearned on van Gogh art style, and investigate four different initialization schemes: Prompt embeddings initialized "close" to the reference target of the erased concept by padding and replacing random tokens ($\pm$ tokens) or characters ($\pm$ characters) of random lengths; Prompt embeddings initialized "far" from the reference target by uniformly sampling random tokens or characters. Remarkably, embeddings initialized in all four regions of the embedding space can successfully restore the erased concept with short optimization trajectories (Figure 2a and 2b). In particular, embeddings initialized in the far region can generate the erased concept without approaching the region close to the reference target embedding. This suggests that, there often locally exists an adversarial prompt embedding close to the initialization. In other words, unlearned models are generally vulnerable to small perturbations to the text embedding, and effective adversarial embeddings are widespread in the prompt embedding space. This echoes the well-established understanding that neural networks are adversarially vulnerable to small perturbation to its inputs Fawzi et al. (2018); Simon-Gabriel et al. (2018); Wu et al. (2020); Beerens & Higham (2024).
For denoiser-based erasure methods like ESD Gandikota et al. (2023a), embeddings initialized "close" to the target embedding tend to diverge from the reference target embedding during optimization (Figure 2c and 2d). This suggests that denoiser-based erasure methods only suppress generation of a concept in a localized region around the reference target embedding. By contrast, for methods that unlearn the text encoders, such as AdvUnlearn Zhang et al. (2024b), embeddings tend to converge slightly towards yet still remain far away from the reference target. This dynamic indicates a different failure mode: modifying the text encoder alters the token-embedding mapping but does not necessarily erase the model’s inherent ability to generate the concept from embeddings located in that specific region.
Collectively, these results demonstrate that embedding-level vulnerabilities are ubiquitous and remain largely unaddressed except in some specific regions of the embedding space. The pervasive nature of these embeddings make the concept-erased models susceptible to exploitation by token-level restoration algorithms. Conversely, for erasure methods that are more robust to embedding-level attacks, such as SH Wu & Harandi (2024) (see Appendix B), token-level attacks are also likely to underperform. Building on this, we further elaborate in Appendix A that the pervasiveness of these vulnerabilities is not an artifact of the concept erasure process, but inherited from the original pre-unlearned model: it is possible to find local embeddings for generating a target concept even when initializing from a semantically distant starting point.
3.3 Token-Level Attacks: New Method
Existing concept restoration methods use gradient-based optimization, which necessitates the projection of the non-differentiable discrete text prompts to a continuous and differentiable space Chin et al. (2023); Zhang et al. (2023). However, recent studies have observed superior performance of coordinate-descend-based methods over their projection counterparts when optimizing on discrete texts Carlini et al. (2023); Zou et al. (2023); Jones et al. (2023). This motivates us to introduce RECORD, a coordinate descent approach that iteratively optimizes the prompt by refining one token at a time while fixing all other tokens. A naive implementation of this strategy requires evaluating the loss function for every token in the vocabulary at each position, which quickly becomes intractable for large vocabularies. To make this optimization feasible, RECORD uses a two-stage approach of leveraging a linear approximation of the loss gradient to identify a small subset of candidate tokens, then perform exact evaluations to determine the optimal token for substitution.
More precisely, the algorithm first initializes a random token sequence $y=[y_{1},\ldots,y_{S}]$ of length $S$, and iteratively performs $N$ passes over $y$. In each pass, a random permutation of the token positions is generated to mitigate positional bias during updates. For each position $s$ in the permuted sequence, the algorithm samples a batch of clean latents $z_{0}^{[n]}$ and corresponding timesteps $t^{[n,s]}$. Candidate tokens $v$ for position $s$ are then selected by sampling $J$ random tokens, and computing the gradient $g$ of the loss $\hat{L}$ from Equation [2] with respect to the candidate token embeddings:
$g_{j}=\nabla{v_{j}}\hat{L}(\mathcal{T}(y({v_{j}},s)),\mathcal{Z}(n,s)).$